OpenStudy (anonymous):

Simplify as much as possible. -9/4w^3 + 5/2w^2 = Can someone help me solve this

OpenStudy (mathstudent55):

The two terms are not like terms, so they can't be combined.

OpenStudy (mathstudent55):

You can factor out a common factor if that what you are being asked to do.

OpenStudy (anonymous):

The question says simplify as much as possible.

OpenStudy (anonymous):

what do you get when you factor it?

OpenStudy (mathstudent55):

Factor \(\dfrac{1}{2} w^2\) from both terms.

OpenStudy (anonymous):

ok thanks so much!! what did you get as your final answer?

OpenStudy (anonymous):

did you get this? 9w34+5w22

OpenStudy (mathmath333):

\(\large \begin{align} \color{black}{\dfrac{-9}{4w^3 }+ \dfrac{5}{2w^2}\hspace{.33em}\\~\\ =\dfrac{-9}{4w^3 }+ \dfrac{5}{2w^2}\hspace{.33em}\\~\\ \normalsize \text{now we have to take least common denominator} \hspace{.33em}\\~\\ =\dfrac{-9}{4w^3 }+ \dfrac{5\times 2w}{2w^2\times 2w}\hspace{.33em}\\~\\ =\dfrac{-9}{4w^3 }+ \dfrac{10w}{4w^{2+1}}\hspace{.33em}\\~\\ =\dfrac{-9}{4w^3 }+ \dfrac{10w}{4w^{3}}\hspace{.33em}\\~\\ =\dfrac{-9+10w}{4w^3 }\hspace{.33em}\\~\\ }\end{align}\)

OpenStudy (anonymous):

Thank you!!

OpenStudy (mathstudent55):

I thought that w^3 and w^2 were being multiplied by numbers. The way it's written in the original question, w^3 and w^2 are not in the denominators.